find the equation
passing through (1,7) and (3,3)

Answer by Alan3354(69443)

find the equation
passing through (1,7) and (3,3)
---------------
Find the slope m
m = (y2-y1)/(x2-x1)
m = (3-7)/(3-1) = -4/2
m = -2
Use either point
y-y1 = m*(x-x1)
y-7 = -2*(x-1)
y-7 = -2x+2
y = -2x + 9 (slope-intercept form)
2x + y = 9 (standard form)